Calculus II Spring 2002 MA114:004–006 Russell Brown

• Homework #14 §10.4 #1, 3, 5, 7, 9, 11, 33, 35, 42, 43. §10.5 #1, 3, 5, 7, 9, 11, 13, 22, 23, 25, 26, 27.

• Homework #15 §10.6 #1, 3, 5, 7, 9, 11, 15, 29, 35.

• Homework #14 should be discussed on Thursday, 21 March 2002. We will not have gotten very far into alternating series. Finish homework 14 and discuss homework 15 on Tuesday, 26 March 2002.

• The limit comparison test is very easy to use, but it may not be so clear why it works. Thus, you should also know the ordinary comparison test.

• The alternating series test gives examples of series which are conditionally convergent, but not absolutely convergent. As you read section 10.6, be sure you know what these words mean.

• In section 10.6, we will not consider the root test. The ratio test is enough for most applications.

• The alternating series test gives a simple way to estimate the error committed when approximating the series by a partial sum. The integral test also gives a way to estimate the error committed when approximating a series by a partial sum. This might make a good test question.

• Homework H. Due Wednesday, 27 March 2002. Section 10.3 #24. Section 10.6 #28.

• Extra credit opportunity. Attend the Math club lecture on Wednesday, 27 March 2002. Write a brief description of what you learned about the mathematics of soap bubbles and hand it in on Friday, 29 March 2002. You may earn up to 5 homework points.

• Quiz 8. Thursday, 28 March 2002. Sections 10.3-10.6.

• Please be sure you have the correct time for the final exam. Our final will be on Thursday, 2 May 2002 at 8am. The date given on the course calendar is wrong.

March 20, 2002